Suppose Tom has 17 coins totaling $2.60. If he has only dimes and quarters, how many of each type does he have?

Answer by addingup(3677)

d + q = 17 subtract q from both sides:
d = 17 - q We'll use this value for d below:
0.10d + 0.25q = 2.60
0.10(17-q) + 0.25q = 2.60
1.7 - 0.10q + 0.25q = 2.60
0.15q = 0.9
q = 0.9/0.15 = 6 he has 6 quarters
d = 17- 6 = 11 he has 11 dimes
----------------------------------------
check:
6(0.25) + 11(0.10) = 2.60
1.5 + 1.10 = 2.60 Correct